# What’s Mine becomes Yours: Defining, Annotating and Detecting Context-Dependent Paraphrases in News Interview Dialogs

### v1

@@ B.3 Paraphrase Candidate Selection
PARA. After selecting the RANDOM set, the PARA set of 400 (guest, host)-pairs was sampled to reach a specified total 350 paraphrases and 150 non-paraphrases together with the RANDOM set.212121 RANDOM and PARA were undergoing annotation together in a second annotation round, after BALANCED had already been annotated. The aim was to reach a higher distribution of paraphrases in our released dataset. The 350/150 split was somewhat arbitrary. It could have easily been 400/100 or 300/200 as well. The PARA set was selected to make the total number of non-repetition paraphrases together with RANDOM reach 300, while limiting the amount of repetition paraphrases to 50. Conversely, non-paraphrases were sampled to add up to 150. This led to 66 non-paraphrases and 334 paraphrases being sampled for the PARA set.
@@
Self-Disagreement Guest: [..] So there was a consensus organization last year that people from genetics and ethics law got together and said, in theory, it should be acceptable to try this in human beings. The question will be, how much safety and evidence do we have to have from animal models before we say it’s acceptable. Host: When it comes to this issue, let’s face it, while there are the concerns here in the United States, it’s happening in other countries.
(C) distinguish paraphrases from inferences, conclusions or “just” highly related utterances
@@
Guest: Yes, it is, all \$640,000. Host: That’s a lot of dough.
(CD) emphasize situational aspect to annotators, (H) ask for token-level accuracy of highlights
Prolific Annotators Guest: […] He was an employee that worked downtown Cleveland and saw it fall out of the armored car carrier, and pick it up, and took it, and placed it in his car. Host: And he’s been holding it ever since?
similar to (C)
Guest: […] Would I ever thought that this would be happening, no, it is, it’s crazy? Just enjoy the moment. Host: […] , Magic Johnson was saying that when he first started taking meetings with investors or with business people, they didn’t take him seriously, but he thought maybe they just wanted his autograph. […]
(AT) use annotator screening to throw out annotators more likely to produce non-sensical pairs
Guest: […] they say, you, you must sue “Fortnite”, and I’m like, “Fortnite”, what is that? I don’t even know what it is Host: So you weren’t even familiar?
(AT) throw out annotators that do not select obvious pairs
Table 14: Examples of Disagreements in Paraphrase Annotation Pilots. All of the presented examples were highlighted by at least one annotator and selected as not showing any paraphrases at all by at least one other annotator. We show examples from three different conditions: Self-disagreement for the lead author, disagreements between volunteers/lab members and disagreements between Prolific annotators. These disagreements informed later training instructions: For (C), see Figure 6; for (P), see Figure 9; for (CD), see Figure 10; for (H), see Figure 8; for (AT), we chose the separate training setup with attention and comprehension checks, see Figures 5, 11 and 13. Early on, we chose to include repetitions in our paraphrase definition since it turned out to be conceptually difficult to separate the two – especially in a context-dependent setting (e.g., is “You don’t know.” a repetition of “I do not know it.” or not?), see Figure 4.
@@ Appendix C Annotations
We train participants to recognize paraphrases (see Figure 412 for the instructions they received). We presented (guest, host)-pairs with their MediaSum summaries, the date of the interview and the interviewer names for context.222222 The additional information of summary, date and speaker names increased reported understanding of context and eased difficulty of the task in pilot studies among lab members. Participants were only admitted to the paraphrase annotation if they passed two attention checks (see Figure 13) and two comprehension checks (see Figure 5 and 11).
@@
Figure 11: Annotator Training (8). Example of an accepted answer for the comprehension check at the end. Only annotators selecting a version of this are admitted to annotate unseen instances. Some of the admitted annotators additionally selected the pair “he’s improved a lot” and “he’s expected to make a full recovery”.
@@
Figure 15: Annotator Recruitment Strategies. To decide the number of annotators for a specific item, we test three different strategies: (1) using a fixed number of annotators across all items (ALL), (2) increasing the number of annotators until at least $n$ annotators agree for each item (absolute) and (3) increasing the number of annotators from 3 until the entropy is smaller than a given threshold (entropy) or a maximum of 10, 15 or 20 annotators is reached. We display the accuracy of the methods compared to using all 20 annotations in (14(a)) and the reliability measure kRR depending on the average number of annotators used (Wong and Paritosh, 2022) in (14(b)). We set a maximum average cost of 8 annotators per item and require a minimum accuracy of 90% as well as a minimum kRR of 0.70. When a strategy fulfills these requirements (i.e., falls in the upper left quadrants for (a) and (b)), we display the entropy thresholds for (3) and absolute number of annotators for (2).
@@ C.4 Annotator Allocation Strategy
Via Prolific’s internal screening system, we recruited native speakers located in the US. Payment for a survey was only withheld if annotators failed two attention checks within the same survey or when a comprehension check at the very beginning of the study was failed [^24] Technically, in line with Prolific guidelines, we do not withhold payment but ask annotators to “return” their study in this case. Practically this is the same, as all annotators did return such a study when asked. in line with Prolific guidelines. [^25] Prolific Attention and Comprehension Check Policy Across all Prolific studies performed for this work (including pilots), we payed participants a median of $8.98\pounds/h\approx 11.41\$/h$262626on March 20th 2024 which is above federal minimum wage in the US. [^27] Federal minimum wage in the US is \$$\$7.25/h\approx 5.71\pounds/h$ according to https://www.dol.gov/agencies/whd/minimum-wage on March 20th 2024
@@
Figure 16: On BALANCED, later training sessions take longer and pass fewer quality checks. In 15(a), we display the seconds the nth annotator needs to go through the training session. The annotators are ordered according to the dates they completed training. Annotations were distributed across 6 different days in June 2023. The green line represents the median duration time of the first n participants. The red line displays the initially estimated completion time of 900 seconds according to pilot studies. The blue line is a linear regression estimate of the duration and it’s 95% confidence interval. On average, participants participating on a later date need more time to finish. In 15(b), we display the summed number of the first n participants that passed the quality checks during training. The grey line represents the angle bisector, i.e., if every participant would pass all quality checks. Later participants are less likely to pass the quality checks.
@@
C.7 Anonymization
@@ D.1 In-Context Learning
A Paraphrase is a rewording or repetition of content in the guests statement. It rephrases what the guest said. Given an interview on - with the summary: Fresh Prince Star Alfonso Ribeiro Sues Over Dance Moves; Rapper 2 Milly Alleges His Dance Moves were Copied. Guest and Host say the following: Guest (TERRENCE FERGUSON, RAPPER): I guess it was season 5 when they premiered it in the game. A bunch of DMs, a bunch of Twitter requests, e-mails, everything was like, you, your game is in the dance, you need to sue, "Fortnite" stole it. Even like big artists, major artists like Joe Buttons and stuff, they have their own like show, daily struggle, they say, you, you must sue "Fortnite", and Im like, "Fortnite", what is that? I dont even know what it is  Host (QUEST): So you werent even familiar? In the reply, does the host paraphrase something specific the guest says? Explanation: Lets think step by step. Terrence Ferguson says at the end of his turn that he didnt know Fortnite. Quest, the host of the interview, repeats that the guest doesnt know Fortnite. So they both say that the guest didnt know Fortnite. Therefore, the answer is yes, the host is paraphrasing the guest. Verbatim Quote Guest: "Im like, "Fortnite", what is that?  I dont even know what it is" Verbatim Quote Host: "you werent even familiar?" Classification: Yes. Given an interview on 2013-10-1 with the summary:  Guest and Host say the following: Guest (REP. RAUL LABRADOR (R), IDAHO):  Host (BLITZER):  In the reply, does the host paraphrase something specific the guest says? Explanation: Lets think step by step. EXPLANATION Therefore, the answer is yes, host is paraphrasing the guest. Verbatim Quote Guest: "We would like the senators to actually come and negotiate with us." Verbatim Quote Host: "you want to negotiate" Classification: Yes. ITEM Explanation:  Verbatim Quote Guest: None. Verbatim Quote Host: None. Classification: No. ITEM Explanation:  Verbatim Quote Guest: "She" "Talked about family life." "errands they need to run and things like that." Verbatim Quote Host: "she talked" "about her family and her kids." "how theyre living day by day." Classification: Yes. ITEM Explanation:  Verbatim Quote Guest: None. Verbatim Quote Host: None. Classification: No. ITEM Explanation:  Verbatim Quote Guest: None. Verbatim Quote Host: None. Classification: No. ITEM Explanation:  Verbatim Quote Guest: "shipping him here to me" Verbatim Quote Host: "coming to New Jersey and being under the auspices" "of De Lacy Davis." Classification: Yes. ITEM Explanation:  Verbatim Quote Guest: "Im to see him." Verbatim Quote Host: "him" "have a visit from you" Classification: Yes. Given an interview on DATE with the summary: SUMMARY Guest and Host say the following: Guest (NAME): UTTERANCE Host (NAME): UTTERANCE Explanation: Lets think step by step.
@@
Table 15: Hyperparameter tuning on the DEV set. We train a token classifier for learning rates 1e-3, 3e-3, 5e-3 and epochs 4, 8, 12 and 16 for 3 seeds. We keep learning rate fixed at 3e-3 when varying the number of epochs and epoch fixed at 8 when varyig the learning rates.
@@ D.2 Token Classification
We use settings very close to Wang et al. (2022a) and test different learning rates and number of epochs with 3 different seeds each. We use the "save best model" option to save the model after the epoch which yielded the best result on the dev set. For the results, see Figure 15.
D.3 Error Analysis
See more errors by the models in Table 16.
Hallucinations. One of the biggest problems for in-context learning are the extractions of the highlighting from the model responses which has values up to 55% in Table 7. Most of these errors can be split into two categories: (1) inconsistent highlighting, where the model classifies a paraphrase but does not highlight text spans in both, the guest and host utterance and (2) hallucinations, where the model highlights spans that do not exist in that form in the guest or host utterance. Hallucination is more prevalent than inconsistent highlighting for GPT-4, where in most cases it leaves out words (e.g., “coming back to a normal winter” vs. “coming back daryn to a normal winter”), in some other cases it adds or replaces words (e.g.,“he’s a counterpuncher” vs. “he’s counterpuncher”), uses morphological variation (e.g., “you’ve” vs. “you have”) or quotes from the wrong source (e.g., from the host when considering the guest utterance). When looking at the classifications by GPT-4 they often seem plausible, even when counted as incorrect with the F1 score. However, even when provided, the highlights are sometimes of low quality.
@@ D.4 Computing Infrastructure
Model M C Shortened Examples
DeBERTa G: There are people that are in that age range where we know they’re high risk, why are they going to the supermarket to buy their own groceries? Get the community, the neighborhood to go and help them. H: if you’re going to help somebody by helping them maybe get their groceries, how long does the coronavirus live on surfaces?
GPT-4 G: We also want to see what connections exist between pardons and potential gifts to the Clinton Library. H: Congressman, short of, though, having a thank-you note attached to a check that went to the Clinton Library, what is it exactly that is going to prove that there was a quid pro quo, that these pardons were actually bought?
Table 16: Errors GPT-4 and DeBERTa. Paraphrases predicted by DeBERTa and GPT-4 but not by the crowd. We display the classification by the model (M) and the crowd majority vote (C). If the model did predict a paraphrase the corresponding text spans are underlined. For comparison, we also display the human highlights if they exist. Note, highlights can exist even if the crowd majority vote did not predict a paraphrase.

### v2

@@ B.3 Paraphrase Candidate Selection
PARA. After selecting the RANDOM set, the PARA set of 400 (guest, host)-pairs was sampled to reach a specified total 350 paraphrases and 150 non-paraphrases together with the RANDOM set.282828 RANDOM and PARA were undergoing annotation together in a second annotation round, after BALANCED had already been annotated. The aim was to reach a higher distribution of paraphrases in our released dataset. The 350/150 split was somewhat arbitrary. It could have easily been 400/100 or 300/200 as well. The PARA set was selected to make the total number of non-repetition paraphrases together with RANDOM reach 300, while limiting the amount of repetition paraphrases to 50. Conversely, non-paraphrases were sampled to add up to 150. This led to 66 non-paraphrases and 334 paraphrases being sampled for the PARA set.
@@
Self-Disagreement Guest: [..] So there was a consensus organization last year that people from genetics and ethics law got together and said, in theory, it should be acceptable to try this in human beings. The question will be, how much safety and evidence do we have to have from animal models before we say it’s acceptable. Host: When it comes to this issue, let’s face it, while there are the concerns here in the United States, it’s happening in other countries. (C) distinguish paraphrases from inferences, conclusions or “just” highly related utterances
@@
Guest: Yes, it is, all \$640,000. Host: That’s a lot of dough. (CD) emphasize situational aspect to annotators, (H) ask for token-level accuracy of highlights
Prolific Annotators Guest: […] He was an employee that worked downtown Cleveland and saw it fall out of the armored car carrier, and pick it up, and took it, and placed it in his car. Host: And he’s been holding it ever since? similar to (C)
Guest: […] Would I ever thought that this would be happening, no, it is, it’s crazy? Just enjoy the moment. Host: […] , Magic Johnson was saying that when he first started taking meetings with investors or with business people, they didn’t take him seriously, but he thought maybe they just wanted his autograph. […] (AT) use annotator screening to throw out annotators more likely to produce non-sensical pairs
Guest: […] they say, you, you must sue “Fortnite”, and I’m like, “Fortnite”, what is that? I don’t even know what it is Host: So you weren’t even familiar? (AT) throw out annotators that do not select obvious pairs
Table 15: Examples of Disagreements in Paraphrase Annotation Pilots. All of the presented examples were highlighted by at least one annotator and selected as not showing any paraphrases at all by at least one other annotator. We show examples from three different conditions: Self-disagreement for the lead author, disagreements between volunteers/lab members and disagreements between Prolific annotators. These disagreements informed later training instructions: For (C), see Figure 6; for (P), see Figure 9; for (CD), see Figure 10; for (H), see Figure 8; for (AT), we chose the separate training setup with attention and comprehension checks, see Figures 5, 11 and 12. Early on, we chose to include repetitions in our paraphrase definition since it turned out to be conceptually difficult to separate the two – especially in a context-dependent setting (e.g., is “You don’t know.” a repetition of “I do not know it.” or not?), see Figure 4.
@@ Appendix C Annotations
We train participants to recognize paraphrases (see Figure 413 for the instructions they received). We presented (guest, host)-pairs with their MediaSum summaries, the date of the interview and the interviewer names for context.292929 The additional information of summary, date and speaker names increased reported understanding of context and eased difficulty of the task in pilot studies among lab members. Participants were only admitted to the paraphrase annotation if they passed two attention checks (see Figure 12) and two comprehension checks (see Figure 5 and 11).
@@
Figure 11: Annotator Training (8). Example of an accepted answer for the comprehension check at the end. Only annotators who highlighted similar spans are admitted to annotate unseen instances. Some of the admitted annotators additionally selected the pair “he’s improved a lot” and “he’s expected to make a full recovery”.
@@
Figure 15: Annotator Recruitment Strategies. To decide the number of annotators for a specific item, we test three different strategies: (1) using a fixed number of annotators across all items (ALL), (2) increasing the number of annotators until at least $n$ annotators agree for each item (absolute) and (3) increasing the number of annotators from 3 until the entropy is smaller than a given threshold (entropy) or a maximum of 10, 15 or 20 annotators is reached. We display the accuracy of the methods compared to using all 20 annotations in (15(a)) and the reliability measure kRR depending on the average number of annotators used (Wong and Paritosh, 2022) in (15(b)). We set a maximum average cost of 8 annotators per item and require a minimum accuracy of 90% as well as a minimum kRR of 0.70. When a strategy fulfills these requirements (i.e., falls in the upper left quadrants for (a) and (b)), we display the entropy thresholds for (3) and absolute number of annotators for (2).
@@ C.4 Annotator Allocation Strategy
Via Prolific’s internal screening system, we recruited native speakers located in the US. Payment for a survey was only withheld if annotators failed two attention checks within the same survey or when a comprehension check at the very beginning of the study was failed [^31] Technically, in line with Prolific guidelines, we do not withhold payment but ask annotators to “return” their study in this case. Practically this is the same, as all annotators did return such a study when asked. in line with Prolific guidelines. [^32] Prolific Attention and Comprehension Check Policy Across all Prolific studies performed for this work (including pilots), we payed participants a median of $8.98\pounds/h\approx 11.41\$/h$333333on March 20th 2024 which is above federal minimum wage in the US. [^34] Federal minimum wage in the US is \$$\$7.25/h\approx 5.71\pounds/h$ according to https://www.dol.gov/agencies/whd/minimum-wage on March 20th 2024
@@
Figure 16: On BALANCED, later training sessions take longer and pass fewer quality checks. In 16(a), we display the seconds the nth annotator needs to go through the training session. The annotators are ordered according to the dates they completed training. Annotations were distributed across 6 different days in June 2023. The green line represents the median duration time of the first n participants. The red line displays the initially estimated completion time of 900 seconds according to pilot studies. The blue line is a linear regression estimate of the duration and it’s 95% confidence interval. On average, participants participating on a later date need more time to finish. In 16(b), we display the summed number of the first n participants that passed the quality checks during training. The grey line represents the angle bisector, i.e., if every participant would pass all quality checks. Later participants are less likely to pass the quality checks.
@@
C.7 Intra-Annotator Annotations Quality
We manually randomly sample ten annotators (with anonymized PROLIFIC ids 60, 6, 86, 84, 47, 31, 68, 88, 41, 92) and analyze 42 of their annotatations. Nine annotators consistently provide plausible annotations, while the other annotator chooses “not a paraphrase” a few times too often. We also noticed some other annotator-specific tendencies, for example, one annotator might tend to highlight fewer words, more words or prefer exact lexical matches.
C.8 Anonymization
@@ D.1 In-Context Learning
{mdframed}
A Paraphrase is a rewording or repetition of content in the guests statement. It rephrases what the guest said.
Given an interview on - with the summary: Fresh Prince Star Alfonso Ribeiro Sues Over Dance Moves; Rapper 2 Milly Alleges His Dance Moves were Copied.
Guest and Host say the following:
Guest (TERRENCE FERGUSON, RAPPER): I guess it was season 5 when they premiered it in the game. A bunch of DMs, a bunch of Twitter requests, e-mails, everything was like, you, your game is in the dance, you need to sue, "Fortnite" stole it. Even like big artists, major artists like Joe Buttons and stuff, they have their own like show, daily struggle, they say, you, you must sue "Fortnite", and Im like, "Fortnite", what is that? I dont even know what it is
Host (QUEST): So you werent even familiar?
In the reply, does the host paraphrase something specific the guest says?
Explanation: Lets think step by step.
Terrence Ferguson says at the end of his turn that he didnt know Fortnite.
Quest, the host of the interview, repeats that the guest doesnt know Fortnite.
So they both say that the guest didnt know Fortnite. Therefore, the answer is yes, the host is paraphrasing the guest.
Verbatim Quote Guest: "Im like, "Fortnite", what is that? I dont even know what it is"
Verbatim Quote Host: "you werent even familiar?"
Classification: Yes.
Given an interview on 2013-10-1 with the summary:
Guest and Host say the following:
Guest (REP. RAUL LABRADOR (R), IDAHO):
Host (BLITZER):
In the reply, does the host paraphrase something specific the guest says?
Explanation: Lets think step by step. EXPLANATION Therefore, the answer is yes, host is paraphrasing the guest.
Verbatim Quote Guest: "We would like the senators to actually come and negotiate with us."
Verbatim Quote Host: "you want to negotiate"
Classification: Yes.
ITEM
Explanation:
Verbatim Quote Guest: None.
Verbatim Quote Host: None.
Classification: No.
ITEM
Explanation:
Verbatim Quote Guest: "She" "Talked about family life." "errands they need to run and things like that."
Verbatim Quote Host: "she talked" "about her family and her kids." "how theyre living day by day."
Classification: Yes.
ITEM
Explanation:
Verbatim Quote Guest: None.
Verbatim Quote Host: None.
Classification: No.
ITEM
Explanation:
Verbatim Quote Guest: None.
Verbatim Quote Host: None.
Classification: No.
ITEM
Explanation:
Verbatim Quote Guest: "shipping him here to me"
Verbatim Quote Host: "coming to New Jersey and being under the auspices" "of De Lacy Davis."
Classification: Yes.
ITEM
Explanation:
Verbatim Quote Guest: "Im to see him."
Verbatim Quote Host: "him" "have a visit from you"
Classification: Yes.
Given an interview on DATE with the summary: SUMMARY
Guest and Host say the following:
Guest (NAME): UTTERANCE
Host (NAME): UTTERANCE
Explanation: Lets think step by step.
@@
Table 16: Hyperparameter tuning on the DEV set. We train a token classifier for learning rates 1e-3, 3e-3, 5e-3 and epochs 4, 8, 12 and 16 for 3 seeds. We keep learning rate fixed at 3e-3 when varying the number of epochs and epoch fixed at 8 when varyig the learning rates. Best options of learning rate and epoch are underlined. Best F1 score is boldfaced.
@@ D.2 Token Classification
We use settings very close to Wang et al. (2022a) and test different learning rates and number of epochs with 3 different seeds each. We use the "save best model" option to save the model after the epoch which yielded the best result on the dev set. For the results, see Figure 16. We use a learning rate of 3e-3 and 12 epochs for further modeling.
D.3 Highlighting Analysis
We compare the highlights provided by DeBERTa AGGREGATED353535i.e., seed $202$ with F1 score of $0.76$, precision of $0.72$ and recall of $0.84$, see https://huggingface.co/AnnaWegmann/Highlight-Paraphrases-in-Dialog and DeBERTa ALL363636i.e., seed $201$ with F1 score of $0.72$, precision of $0.84$ and recall of $0.63$, see https://huggingface.co/AnnaWegmann/Highlight-Paraphrases-in-Dialog-ALL on 10 text pairs from the test set that were classified as paraphrases by both models. We provide examples in Table 17. DeBERTa ALL highlights are shorter, often more on point and arguably more consistent than DeBERTa AGGREGATED highlights. We also manually analyzed 10 text pairs from the test set that GPT-4 classified as paraphrases. We provide examples of GPT-4 highlights in Table 18. Generally, they seem of good quality, but have the tendency to span complete sub-sentences, even if not all is relevant.
Hallucinations. One of the biggest problems for in-context learning are the extractions of the highlighting from the model responses which has errors in up to 71% of the cases in Table 8. Most of these errors can be split into two categories: (1) inconsistent highlighting, where the model classifies a paraphrase but does not highlight text spans in both, the guest and host utterance and (2) hallucinations, where the model highlights spans that do not exist in that form in the guest or host utterance. Hallucination is more prevalent than inconsistent highlighting for GPT-4, where in most cases it leaves out words (e.g., “coming back to a normal winter” vs. “coming back daryn to a normal winter”), in some other cases it adds or replaces words (e.g.,“he’s a counterpuncher” vs. “he’s counterpuncher”), uses morphological variation (e.g., “you’ve” vs. “you have”) or quotes from the wrong source (e.g., from the host when considering the guest utterance). Most of these extraction errors seem to be resolvable by humans when looking at them manually, so it might be possible to address them in future work with a more advanced matching algorithm or by querying GPT-4 until one gets a parsable response. When looking at the classifications by GPT-4 they often seem plausible, even when counted as incorrect with the F1 score.
@@ D.4 Computing Infrastructure
AGG ALL C Shortened Examples
G: There are people that are in that age range where we know they’re high risk, why are they going to thesupermarket tobuy their own groceries? Get the community, the neighborhood to go and help them. H: if you’regoing to help somebodyby helping them maybe get their groceries, how long does the coronavirus live on surfaces?
G: And people always prefer, of course, to see the pope as the principal celebrant of the mass. So that’s good. That’ll be tonight. And it will be his 26th mass and it will be the 40th or, rather, the 30th time that this is offered in round the world transmission. And it will be my 20th time in doing it as a television commentator from Rome so. H: Yes, you’ve been doing this for a while now.
G: Well, what happened was we finally waved down a Coast Guard helicopter. And what they were looking for were people with disabilities and medical conditions, which none of us really had. They didn’t lift any of us into the helicopter or anything. What they told us was to basically walkout of our house, up the street, trying tofight against thecurrent that was going theopposite way of where we needed to go. H: So you walked through that current to get to the higher ground or get to a drier spot?
Table 17: DeBERTa ALL vs DeBERTa AGGREGATED highlights. Paraphrase highlights predicted by the best DeBERTa ALL (i.e., seed 201 with F1 score of 0.72) and the best DeBERTa AGG model (i.e., seed 202 F1 score of 0.76, same as in the main paper). Even though DeBERTa AGG gets better F1 scores on classification, the DeBERTa ALL highlights are arguably more on point. For comparison, we also display the human highlights if they exist. Note, highlights can exist even if the crowd majority vote did not predict a paraphrase.
GPT-4 C Shortened Examples
G: We also want to see what connections exist between pardons and potential gifts to the Clinton Library. H: Congressman, short of, though, having a thank-you note attached to a check that went to the Clinton Library, what is it exactly that is going to prove that there was a quid pro quo, that these pardons were actually bought?
G: [Trump] is appointing very young judges. H: […] if you’re 50-plus, you’re probably too old for the Trump Administration to be seriously considered for a district court judgeship.
Table 18: GPT-4 highlights. Paraphrase highlights predicted by GPT-4. For comparison, we also display the human highlights if they exist. Note, highlights can exist even if the crowd majority vote did not predict a paraphrase.